# AI Governance through Markets

### v1

@@ 4. Protocolisaton: Procurement Standards
Current procurement standards for AI systems are evolving to address challenges of safety, transparency, accountability, and security. Governments often rely on private contractors for AI development, necessitating careful planning during procurement to ensure transparency and compliance with norms (Dor and Coglianese, 2021). Guidelines for AI procurement are being developed to support decision-makers in identifying key elements for secure AI systems, considering technical and regulatory environments (Kieseberg et al., 2022). There is a particularly urgent demand for AI-specific public procurement (Hickok, 2022) standards putside of the defence industry. Private organisations are further developing AI procurement standards, but concerns remain regarding liability and the need for internal governance frameworks (Oketunji et al., 2023; McKinsey and Company, 2024).
@@
Pre-market testing standards, such as the FDA’s 510(k) Premarket Notification process, offer a compelling parallel for how large-scale buyers can use their purchasing power to shape the trajectory of AI development. Just as the 510(k) process expedites regulatory approval for medical devices that demonstrates substantial equivalence to another legally U.S. marketed device, procurement standards for AI could require companies to meet predefined risk assessment benchmarks in order to reduce the friction of entering into new lucrative markets and procurement contracts. This would incentivise a minimum level of efficacy and safety while also fostering innovation within a controlled framework. In the downstream, this may harmonise high AI standards across borders as well. Also, pre-market testing standards typically come with public accountability measures, including the obligation to report adverse events or safety concerns post-market. Such transparency creates a feedback loop, encouraging developers to be proactive in mitigating risks while under the scrutiny of investors, regulators, and the public – as discussed further in the section on “AI disclosures.”
The U.S. Department of Defense’s Cybersecurity Maturity Model Certification (CMMC) exemplifies existing technical safety standards for AI, delineating three-tiered cybersecurity requirements—basic protection, Controlled Unclassified Information (CUI) safeguards, and advanced threat mitigation—with mandatory compliance verification. This framework could inform the development of comprehensive AI procurement standards. However, the establishment of universally applicable AI safety and reliability metrics remains a critical research imperative for practical implementation.
@@ 4.1. Case Study: NASA Apollo Mission
NASA’s procurement framework is structured upon the Federal Acquisition Regulation (FAR) and its organisational supplement, the NASA FAR Supplement (NFS). These regulatory mechanisms are specifically designed to address the distinctive requirements of space exploration, encompassing stringent reliability metrics, safety protocols, and performance criteria (McCaffrey, 2021). NASA’s uncompromising adherence to safety protocols, precision requirements, and performance criteria necessitated enhanced capabilities among contractors (Arrilucea and Commission, 2018).
These protocols catalysed technological innovations that would establish industry benchmarks for subsequent decades, encompassing advancements in propulsion systems, microelectronic components, computational software, materials engineering, navigational systems, and solar energy technology—innovations that were subsequently adapted for civilian applications ((1971), NTRS). Contemporary commercial aviation and aerospace enterprises, including SpaceX and The Boeing Company, maintain comparably stringent safety protocols per the NFS procurement standards (Administration, 2017).
@@ 5. Investor behaviour: Due Diligence
Investor due diligence refers to the comprehensive appraisal conducted by investors to assess the potential risks and rewards associated with an investment opportunity. This process involves the systematic evaluation of financial statements, management competency, market conditions, and regulatory compliance (Snow, 2023). For Economists, due diligence plays a pivotal role in addressing information asymmetry—the imbalance of information between corporate insiders and external investors—which can lead to suboptimal investment decisions and market inefficiencies (Novikov et al., 2018). For Investors, the ‘the goal of due diligence is to make the buyer comfortable enough to go through with the deal’ (Snow, 2023).
Investor due diligence mitigates information asymmetry by enabling investors to obtain and analyse information that may not be readily apparent from public disclosures (Daley et al., 2024). This process encourages issuers to maintain accurate and comprehensive records, knowing that sophisticated investors will scrutinise their disclosures. Due diligence thus acts as a self-regulating mechanism that complements formal regulatory frameworks (Daley et al., 2024).
@@
  • Market Due Diligence: Analysing market size, growth potential, competitive landscape, and customer base. With AI, this extends to evaluating the potential disruptive impact of AI technologies on existing markets - harnessing standardised disclosures to compare AI integrations across competing firms (Howson, 2017).
The analysis that arises in due diligence helps in accurately pricing securities, reflecting all available information. Its governance ability lies in being the determinant process by which investors allocate capital. In so doing it incentives management practices and accountability within firms (Schilling‐Vacaflor and Lenschow, 2021) and shapes competitive dynamics between firms. It is important to note that during the due diligence process, buyers are not solely focused on identifying potential risks. Equally important to investors is the evaluation of the enterprise’s strengths and opportunities. Thus due diligence is both a carrot and a stick, rewarding those organisations that present the least risk for investors while promising attractive returns (Hudson and Stavrand, 2007).
Significant uncertainty exists around the financial risk of firms with high levels of AI adoption (Ameye et al., 2023). Due diligence can help scrutinise AI practices, in doing so investors help establish market norms for transparency in AI development. Companies are further incentivized to maintain high standards in operational practices and regulatory compliance, knowing that investors are paying close attention to these aspects during the due diligence process (Charles et al., 2023).
@@ 5.1. Case Study: BP Oil Spill
The 2010 BP Deepwater Horizon incident, resulting in the discharge of approximately 4.9 million barrels of oil into the Gulf of Mexico, represents the most extensive marine oil spill in recorded history (Pallardy, [n. d.]). Beyond the catastrophic environmental and economic implications, BP incurred \$65 billion in remediation costs and legal settlements (United States Environmental Protection Agency, 2024). The financial impact was equally significant: BP’s market capitalization decreased by \$105 billion, representing a 55% decline in equity value (Hargreaves, 2010).
@@ 7. Disclosing AI’s development and deployment
Most non-financial disclosures are not obligatory, yet 1,395 investors, who collectively manage \$59 trillion—accounting for 60% of global assets under management—have endorsed the UN Principles of Responsible Investment (Association, 2024). This endorsement signifies a commitment to, among other considerations, take non-financial disclosures into account when making investment decisions. Some funds will include or exclude potential investments on the basis of their non-financial information thus creating a strong incetive (Curtis et al., 2021; Dikolli et al., 2022). There is also large investor demand for non-financial disclosure(Thewissen et al., 2024). For example, it is common for investors to prefer funds that do not invest in companies providing services to the military (Starr, 2008). Instances of widespread, non-financial disclosures encompass research and development (R&D) information (Koh and Reeb, 2014), sustainability reports (Saini et al., 2022), and organisational practices (Arvidsson, 2011). Crucial for our purposes, non-financial disclosures provide market actors with information on potential risks that may not be immediately apparent from financial statements (Pozzoli and Raffaele, 2022; Flöstrand and Ström, 2006).
@@ 7.1. AI Disclosures
We argue that the AI research communities can contribute to AI governance by standardising the disclosure of AI. This standardisation requires the consistent description and quantification of key aspects of AI systems - to enable firms, policymakers and other actors to make informed decisions.
Providers, such as developers of foundational general-purpose AI systems like GPT-4o, create the underlying models, while distributors use these models to offer products, such as ChatGPT (Gutierrez et al., 2023). A growing community of researchers in academia, civil society, government, and industry are actively studying the risks associated with AI. These risks are not always negative, but rather descriptions of various features of the AI models and organisations that develop or distribute them that are currently uncertain. These risks are increasingly communicated through various market governance mechanisms. For instance, organisations like METR, which function as safety auditors, evaluate the risk profiles of companies like Anthropic or OpenAI (Achiam et al., 2023). Over this decade, the standardisation of how AI-related information is transmitted through markets will likely solidify, much as we already see with performance benchmarks such as GPQA (Rein et al., 2023), MATH (Hendrycks et al., 2021), and MMLU (Hendrycks et al., 2020) being widely adopted by companies like OpenAI, DeepMind, and Anthropic.
@@
By establishing clear guidelines for describing risks— such as the allocation of compute, incident reporting, and capabilities in various areas—organisations can provide a comprehensive view of their AI models’ vulnerabilities and limitations. This would be paired with specific metrics to quantify these risks, such as model error rates, data privacy breach incidents, or energy consumption during model training and deployment. Such quantification enables stakeholders to make more informed comparisons between different AI systems and providers. With standardised reporting, market actors can assess AI risks more accurately, reducing uncertainty and promoting better capital allocation.
The aforementioned market governance mechanisms all require a framework of AI disclosure that identifies different sources of risk. To disclose requires an understanding of the different areas that are present in the development and deployment of AI, as well as in the organisations that develop and deploy. For example, OpenAI’s preparedness framework identifies categories such as: 1) Individualised persuasion, 2) Cybersecurity, 3) Chemical, biological, radiological, and nuclear threats, and Autonomous replication and adaptation (OpenAI, 2023). Many other such frameworks exist, with the AI Risk Repository capturing 700+ risks extracted from 43 existing frameworks (Slattery et al., 2024). As such, this present paper will not be providing an exhaustive literature review of different AI Risks.
@@ 8.1. Intellectual Property and Disclosure: A Tradeoff
Artificial intelligence enterprises maintain substantial opacity regarding their operational processes, a phenomenon particularly pronounced in the semiconductor sector where competitive pressures and industrial espionage present significant concerns (Constantin, 2024). However, empirical evidence suggests the possibility of achieving equilibrium between transparency and proprietary protection. Patent systems exemplify such balanced disclosure mechanisms (Institution, 2011). Standardised globally, patent legislation mandates comprehensive technical documentation of innovations, ensuring replicability by qualified practitioners, as codified in Article 83 of the European Patent Convention. This reciprocal arrangement confers temporary market exclusivity in exchange for public knowledge dissemination.
The pharmaceutical industry provides an instructive paradigm for managing the tension between disclosure requirements and intellectual property protection. Global pharmaceutical research and development expenditure exceeded \$300 billion in 2023, representing significant growth from \$5 billion in 1980 and \$38 billion in 2000 (Report, 2024c). Leading pharmaceutical corporations, including Merck, Novartis, and Roche, maintain substantial R&D investment ratios ranging from 27% to 50% of revenue (Report, 2024b). Their disclosure protocols prioritise methodological transparency over proprietary formulations (Deloitte, 2024b), encompassing clinical trial methodologies, regulatory milestones, development timelines, R&D intensity metrics, and scientific uncertainties, while maintaining competitive advantages through strategic patent protection (on Corporate Governance, 2021). This disclosure framework enables investor assessment of future profitability while establishing realistic expectations. The model has facilitated substantial capital attraction despite inherent developmental risks, supported by the potential for significant long-term revenue from successful therapeutic interventions.
The pharmaceutical sector’s approach offers valuable insights for AI disclosure frameworks (to Medicine Foundation, 2020). Strategic transparency—encompassing risk categorization, regulatory compliance milestones, and mitigation protocols—can be achieved without compromising proprietary algorithms. This methodology, focusing on process disclosure rather than technical specifications, effectively balances regulatory compliance, investment attraction, and stakeholder trust. Thus, AI disclosure requirements may be reconceptualized as strategic opportunities rather than regulatory burdens.
@@ 8.2. Market incentives for risk mitigation can unlock funding for safety research
This potential capital reallocation can be contextualised against ESG mandates, where projections indicate asset managers will oversee \$33.9 trillion in ESG-focused funds by 2026 (Report, 2024a). This precedent demonstrates institutional investors’ capacity to integrate non-traditional criteria into investment decisions, despite uncertainty regarding ESG’s correlation with financial returns. Crucially, the primary critique of ESG funds lies in their inconsistent metrics and lack of standardization, which undermine comparability and transparency across firms and industries. This results in fostering superficial compliance practices, such as greenwashing, rather than driving substantive improvements in sustainability or ethical governance. Our argument for a standardized disclosure framework seeks to avoid an equivalent ”safety washing”. Further, the more direct financial implications of AI-related risks may present even more compelling incentives for institutional investors. Thus, strategic influence on capital flows could create substantial financial incentives for responsible AI development on a global scale.
@@ 8.3. Scala Politica: Tailoring Intervention to Size without stifling innovation
The interconnected nature of market actors can create cascading effects throughout the market ecosystem. Institutional investors, particularly private equity firms managing pension fund portfolios, demonstrate how modifications in compliance expectations at upper market echelons can fundamentally reshape capital allocation patterns throughout the investment hierarchy. The dramatic 473.5% increase in Fortune 500 companies identifying AI-related risks underscores this growing concern (Fortune, 2024). Conservative projections suggest that AI-related incidents resulting in merely a 2% devaluation—whether through market capitalization decline, regulatory penalties, or customer attrition—could precipitate substantial financial implications, given the magnitude of capital exposed.
For startups, however, the landscape is starkly different. The financial scale disparity between Fortune 500 corporations and venture capital markets is substantial, with the former generating collective revenues exceeding \$18 trillion in 2023, while venture capital investments totalled approximately \$394 billion in 2022, declining from a peak of \$611 billion in 2021 (Institute, 2023). Due to the complexities inherent in AI supply chains, there are financial interdependencies between emergent enterprises and established corporations (Cahn, 2024). Major corporations, notably cloud service providers, often mitigate the risks startups face through substantial purchase commitments and preferential pricing strategies. While this arrangement currently maintains supply chain stability, its sustainability remains contingent upon the continued risk-absorption practices of technology conglomerates.
@@
The imposition of extensive disclosure requirements on early-stage enterprises may impede innovation trajectories. Resource constraints can limit these organizations’ capacity to navigate these frameworks, resulting in increased operational costs that may deter market entry and extend commercialization timelines (Van Reenen et al., 2021). Excessive regulatory oversight can foster risk aversion, potentially inhibiting the experimental approaches essential for technological advancement (Magazine, 2024). Thus the appropriate level of disclosure must be adjusted to an entity’s size.

### v2

@@ 4. Protocolisaton: Procurement Standards
Current procurement standards for AI systems are evolving to address challenges of safety, transparency, accountability, and security. Governments often rely on private contractors for AI development, necessitating careful planning during procurement to ensure transparency and compliance with norms (Dor and Coglianese, 2021). Guidelines for AI procurement are being developed to support decision-makers in identifying key elements for secure AI systems, considering technical and regulatory environments (Kieseberg et al., 2022). There is a particularly urgent demand for AI-specific public procurement standards outside of the defence industry(Hickok, 2022). Private organisations are further developing AI procurement standards, but concerns remain regarding liability and the need for internal governance frameworks (Oketunji et al., 2023; McKinsey and Company, 2024).
@@
Pre-market testing standards, such as the FDA’s 510(k) Premarket Notification process, offer a compelling parallel for how large-scale buyers can use their purchasing power to shape the trajectory of AI development. Just as the 510(k) process expedites regulatory approval for medical devices that demonstrates substantial equivalence to another legally U.S. marketed device, procurement standards for AI could require companies to meet predefined risk assessment benchmarks in order to reduce the friction of entering into new lucrative markets and procurement contracts. This would incentivise a minimum level of efficacy and safety while also fostering innovation within a controlled framework. In the downstream, this may harmonise high AI standards across borders as well. In addition, pre-market testing standards typically come with public accountability measures, including the obligation to report adverse events or safety concerns post-market. Such transparency creates a feedback loop, encouraging developers to be proactive in mitigating risks while under the scrutiny of investors, regulators, and the public – as discussed further in the section on “AI disclosures.”
The U.S. The Department of Defense’s Cybersecurity Maturity Model Certification (CMMC) exemplifies existing technical safety standards for AI, delineating three-tiered cybersecurity requirements—basic protection, Controlled Unclassified Information (CUI) safeguards, and advanced threat mitigation—with mandatory compliance verification. This framework can inform the development of comprehensive AI procurement standards. However, the establishment of universally applicable AI safety and reliability metrics is imperative for practical implementation.
@@ 4.1. Case Study: NASA Apollo Mission
NASA’s procurement framework is structured upon the Federal Acquisition Regulation (FAR) and its organisational supplement, the NASA FAR Supplement (NFS). These regulatory mechanisms are specifically designed to address the distinctive requirements of space exploration, encompassing stringent reliability metrics, safety protocols, and performance criteria (McCaffrey, 2021). NASA’s uncompromising adherence to safety protocols, precision requirements, and performance criteria necessitates enhanced capabilities among contractors (Arrilucea and Commission, 2018).
These protocols catalysed technological innovations that would establish industry benchmarks for subsequent decades, encompassing advancements in propulsion systems, microelectronic components, computational software, materials engineering, navigational systems, and solar energy technology—innovations that were subsequently adapted for civilian applications ((1971), NTRS). Contemporary commercial aviation and aerospace enterprises, including SpaceX and the Boeing Company, maintain comparably stringent safety protocols per NFS procurement standards (Administration, 2017).
@@ 5. Investor behaviour: Due Diligence
Investor due diligence refers to a comprehensive appraisal conducted by investors to assess the potential risks and rewards associated with an investment opportunity. This process involves a systematic evaluation of financial statements, management competency, market conditions, and regulatory compliance (Snow, 2023). For economists, due diligence plays a pivotal role in addressing information asymmetry—the imbalance of information between corporate insiders and external investors—which can lead to suboptimal investment decisions and market inefficiencies (Novikov et al., 2018). For investors, the ‘the goal of due diligence is to make the buyer comfortable enough to go through with the deal’ (Snow, 2023).
Investor due diligence mitigates information asymmetry by enabling investors to obtain and analyse information that may not be readily apparent from public disclosures (Daley et al., 2024). This process encourages issuers to maintain accurate and comprehensive records, knowing that sophisticated investors will scrutinise their disclosures. Thus, due diligence acts as a self-regulating mechanism that complements formal regulatory frameworks (Daley et al., 2024).
@@
  • Market Due Diligence: Analysing market size, growth potential, competitive landscape, and customer base. With AI, this extends to evaluating the potential disruptive impact of AI technologies on existing markets, harnessing standardised disclosures to compare AI integrations across competing firms (Howson, 2017).
The analysis that arises in due diligence helps in accurately pricing securities, reflecting all available information. Its governance ability lies in being the determinant process by which investors allocate capital. In so doing, it incentivises management practices and accountability within firms (Schilling‐Vacaflor and Lenschow, 2021) and shapes competitive dynamics between firms. It is important to note that during the due diligence process, buyers are not solely focused on identifying potential risks. Equally important to investors is the evaluation of the enterprise’s strengths and opportunities. Thus, due diligence is both a carrot and a stick, rewarding those organisations that present the least risk for investors while promising attractive returns (Hudson and Stavrand, 2007).
Significant uncertainty exists regarding around the financial risk of firms with high levels of AI adoption (Ameye et al., 2023). Due diligence can help scrutinise AI practices, in doing so investors help establish market norms for transparency in AI development. Companies are further incentivised to maintain high standards in operational practices and regulatory compliance, knowing that investors pay close attention to these aspects during the due diligence process (Charles et al., 2023).
@@ 5.1. Case Study: BP Oil Spill
The 2010 BP Deepwater Horizon incident, resulting in the discharge of approximately 4.9 million barrels of oil into the Gulf of Mexico, represents the most extensive marine oil spill in recorded history (Pallardy, [n. d.]). Beythe catastrophic environmental and economic implications, BP incurred \$65 billion in remediation costs and legal settlements (United States Environmental Protection Agency, 2024). The financial impact was equally significant: BP’s market capitalisation decreased by \$105 billion, representing a 55% decline in equity value (Hargreaves, 2010).
@@ 7. Disclosing AI’s development and deployment
Most non-financial disclosures are not obligatory, yet 1,395 investors, who collectively manage \$59 trillion—accounting for 60% of global assets under management—have endorsed the UN Principles of Responsible Investment (Association, 2024). This endorsement signifies a commitment to, among other considerations, take non-financial disclosures into account when making investment decisions. Some funds will include or exclude potential investments on the basis of their non-financial information thus creating a strong incentive (Curtis et al., 2021; Dikolli et al., 2022). There is also large investor demand for non-financial disclosure (Thewissen et al., 2024). For example, it is common for investors to prefer funds that do not invest in companies providing services to the military (Starr, 2008). Instances of widespread, non-financial disclosures encompass research and development (R&D) information (Koh and Reeb, 2014), sustainability reports (Saini et al., 2022), and organisational practices (Arvidsson, 2011). Crucial for our purposes, nonfinancial disclosures provide market actors with information on potential risks that may not be immediately apparent in financial statements (Pozzoli and Raffaele, 2022; Flöstrand and Ström, 2006).
@@ 7.1. AI Disclosures
We argue that the AI research communities can contribute to AI governance by standardising the disclosure of AI. This standardisation requires the consistent description and quantification of key aspects of AI systems to enable firms, policymakers, and other actors to make informed decisions.
Providers, such as developers of foundational general-purpose AI systems like GPT-4o, create the underlying models, while distributors use these models to offer products, such as ChatGPT (Gutierrez et al., 2023). A growing community of researchers in academia, civil society, government, and industry is actively studying the risks associated with AI. These risks are not always negative, but rather descriptions of various features of the AI models and organisations that develop or distribute them that are currently uncertain. These risks are increasingly communicated through various market governance mechanisms. For instance, organisations such as METR, which function as safety auditors, evaluate the risk profiles of companies such as Anthropic or OpenAI (Achiam et al., 2023). Over this decade, the standardisation of how AI-related information is transmitted through markets will likely solidify, much as we already see with performance benchmarks such as GPQA (Rein et al., 2023), MATH (Hendrycks et al., 2021), and MMLU (Hendrycks et al., 2020) being widely adopted by companies like OpenAI, DeepMind, and Anthropic.
@@
By establishing clear guidelines for describing risks— such as the allocation of compute, incident reporting, and capabilities in various areas—organisations can provide a comprehensive view of their AI models’ vulnerabilities and limitations. This would be paired with specific metrics to quantify these risks, such as model error rates, data privacy breach incidents, or energy consumption during model training and deployment. Such quantification enables stakeholders to make informed comparisons between different AI systems and providers. With standardised reporting, market actors can assess AI risks more accurately, reducing uncertainty and promoting better capital allocation.
The aforementioned market governance mechanisms require a framework for AI disclosure that identifies different sources of risk. To disclose requires an understanding of the different areas that are present in the development and deployment of AI, as well as in the organisations that develop and deploy. For example, OpenAI’s preparedness framework identifies categories such as: 1) Individualised persuasion, 2) Cybersecurity, 3) Chemical, biological, radiological, and nuclear threats, and Autonomous replication and adaptation (OpenAI, 2023). Many other such frameworks exist, with the AI Risk Repository capturing 700+ risks extracted from 43 existing frameworks (Slattery et al., 2024). As such, this present paper will not be providing an exhaustive literature review of different AI Risks.
@@ 8.1. Intellectual Property and Disclosure: A Tradeoff
Artificial intelligence enterprises maintain substantial opacity regarding their operational processes, a phenomenon particularly pronounced in the semiconductor sector where competitive pressures and industrial espionage present significant concerns (Constantin, 2024). However, empirical evidence suggests the possibility of achieving equilibrium between transparency and proprietary protection. Patent systems exemplify such balanced disclosure mechanisms (Institution, 2011). Standardised globally, patent legislation mandates comprehensive technical documentation of innovations, ensuring replicability by qualified practitioners, as codified in Article 83 of the European Patent Convention. This reciprocal arrangement confers temporary market exclusivity in exchange for the dissemination of public knowledge.
The pharmaceutical industry provides an instructive paradigm for managing the tension between disclosure requirements and intellectual property protection. Global pharmaceutical research and development expenditure exceeded \$300 billion in 2023, representing significant growth from \$5 billion in 1980 and \$38 billion in 2000 (Report, 2024c). Leading pharmaceutical corporations, including Merck, Novartis, and Roche, maintain substantial R&D investment ratios, ranging from 27% to 50% of revenue (Report, 2024b). Their disclosure protocols prioritise methodological transparency over proprietary formulations (Deloitte, 2024b), encompassing clinical trial methodologies, regulatory milestones, development timelines, R&D intensity metrics, and scientific uncertainties, while maintaining competitive advantages through strategic patent protection (on Corporate Governance, 2021). This disclosure framework enables investors to assess future profitability while establishing realistic expectations. The model has facilitated substantial capital attraction despite inherent developmental risks, supported by the potential for significant long-term revenue from successful therapeutic interventions.
The pharmaceutical sector’s approach offers valuable insights for AI disclosure frameworks (to Medicine Foundation, 2020). Strategic transparency—encompassing risk categorization, regulatory compliance milestones, and mitigation protocols—can be achieved without compromising proprietary algorithms. This methodology, focusing on process disclosure rather than technical specifications, effectively balances regulatory compliance, investment attraction, and stakeholder trust. Thus, AI disclosure requirements may be reconceptualised as strategic opportunities rather than regulatory burdens.
@@ 8.2. Market incentives for risk mitigation can unlock funding for safety research
This potential capital reallocation can be contextualised against ESG mandates, where projections indicate asset managers will oversee \$33.9 trillion in ESG-focused funds by 2026 (Report, 2024a). This precedent demonstrates institutional investors’ capacity to integrate non-traditional criteria into investment decisions, despite uncertainty regarding ESG’s correlation with financial returns. Crucially, the primary critique of ESG funds lies in their inconsistent metrics and lack of standardisation, which undermines comparability and transparency across firms and industries. This results in fostering superficial compliance practices, such as greenwashing, rather than driving substantive improvements in sustainability or ethical governance. Our argument for a standardised disclosure framework seeks to avoid an equivalent ”safety washing”. Further, the more direct financial implications of AI-related risks may present even more compelling incentives for institutional investors. Thus, strategic influence on capital flows could create substantial financial incentives for responsible AI development on a global scale.
@@ 8.3. Scala Politica: Tailoring Intervention to Size without stifling innovation
The interconnected nature of market actors can create cascading effects throughout the market ecosystem. Institutional investors, particularly private equity firms managing pension fund portfolios, demonstrate how modifications in compliance expectations at upper market echelons can fundamentally reshape capital allocation patterns throughout the investment hierarchy. The dramatic 473.5% increase in Fortune 500 companies identifying AI-related risks underscores this growing concern (Fortune, 2024). Conservative projections suggest that AI-related incidents resulting in merely a 2% devaluation—whether through market capitalisation decline, regulatory penalties, or customer attrition—could precipitate substantial financial implications, given the magnitude of capital exposed.
However, startups have a starkly different landscape. The financial scale disparity between Fortune 500 corporations and venture capital markets is substantial, with the former generating collective revenues exceeding \$18 trillion in 2023, while venture capital investments totalled approximately \$394 billion in 2022, declining from a peak of \$611 billion in 2021 (Institute, 2023). Owing to the complexities inherent in AI supply chains, there are financial interdependencies between emergent enterprises and established corporations (Cahn, 2024). Major corporations, notably cloud service providers, often mitigate the risks that startups face through substantial purchase commitments and preferential pricing strategies. While this arrangement currently maintains supply chain stability, its sustainability remains contingent upon the continued risk-absorption practices of technology conglomerates.
@@
The imposition of extensive disclosure requirements on early-stage enterprises may impede innovation trajectories. Resource constraints can limit these organisations’ capacity to navigate these frameworks, resulting in increased operational costs that may deter market entry and extend commercialisation timelines (Van Reenen et al., 2021). Excessive regulatory oversight can foster risk aversion, potentially inhibiting experimental approaches essential for technological advancement (Magazine, 2024). Thus the appropriate level of disclosure must be adjusted to an entity’s size.